a bus starts its journey from rest with constant acceleration 5m/s^2 find its velocity at t=2sec.and draw v-t graph.
1
Expert's answer
2020-09-14T10:20:09-0400

"v(t) = a\cdot t+v(0)"


"acceleration =5( \\tfrac{m}{s^{2}}\t)"


"v(0)=0" (at rest)


"v(t)=5\cdot t+0"


"v(t)=5t"


"v(2sec)=5\\cdot 2=10( \\tfrac{m}{s}\t)"


velocity - time graph: